OpenStudy (anonymous):

simplify the sum. 1.(8u^3+6u^2+8)+(3u^3-6u+8) simplify the defference. 1. (7w^2-7w-3)-(2w^2+5-2) 2. (8w^2-6w-7)-(3w^2+3w-4) find the GCF of the terms of the polynomial. 1.18x^3+24x^2 CAND SOMEBODY HELP ME PLAESE.!!!!!??//??/

OpenStudy (espex):

What is it you do not understand? For the first three problems you need only combine like terms and for the last one you find the greatest value they have in common.

OpenStudy (anonymous):

imm doin work i have never seen until yesterday lol.

OpenStudy (espex):

I'm sure you understand what "like terms" are, so looking at the problems one at a time, combine all of the like terms, once that is accomplished it will be in its simplified form.

OpenStudy (anonymous):

thats it???? okay thanks:)

OpenStudy (espex):

:) You're welcome.

OpenStudy (anonymous):

For the GCF, you want the expression that divides both (which is "in" both). You can do it in bits, GCF of 18 and 24? (The biggest number that divides both)

OpenStudy (anonymous):

thanks.

OpenStudy (anonymous):

Come on, tell me what you get...

OpenStudy (anonymous):

24???????????????:(

OpenStudy (espex):

What if you think of it like this: 18x^3 = 2*3*3 x*x*x

OpenStudy (anonymous):

immm sorrry! imm very dumb in math.

OpenStudy (anonymous):

Well, you can't divide 24 into 18 (you can but the answer wont be an integer). A systematic way to find the GCF is to factor the numbers like this 24/2 = 12, 12/2 = 6, 6/2 = 3 so 24 = 2*2*2*3

OpenStudy (anonymous):

Ah, EspeX says the same....

OpenStudy (espex):

@estudier has this well in hand. :)

OpenStudy (anonymous):

eSpeX

OpenStudy (anonymous):

So 24 = 2*2*2*3 and 18 = 2*3*3 To get the GCF you want to "take out" the bits that are in both. Can you see how?

OpenStudy (anonymous):

okay i see what you are saying now.

OpenStudy (anonymous):

So for instance there are two 3's in 18 but only one in 24 so you can take out one of them.

OpenStudy (anonymous):

okay.

OpenStudy (anonymous):

What about the 2's?

OpenStudy (anonymous):

r u still there?

OpenStudy (anonymous):

I guess not.

OpenStudy (anonymous):

is the answer to the problem 6?

OpenStudy (anonymous):

lol imm still here i just had to run to the restroom.

OpenStudy (anonymous):

That's the correct answer to the GCF of 24 and 18. Now the next bit, we sort of do it the same. What's the bit of x^3 and x^2 that is "in" both?

OpenStudy (anonymous):

remember x^3 is x*x*x and x^2 = x*x

OpenStudy (anonymous):

are you giving me another math problem to work? lol

OpenStudy (anonymous):

It's just the same only with x's instead of 2's or 3's We are still on your original GCF question.

OpenStudy (anonymous):

id get it?? write it out for me

OpenStudy (anonymous):

Write what ?

OpenStudy (anonymous):

what you are saying

OpenStudy (anonymous):

I thought that's what I was doing....?

OpenStudy (anonymous):

lol seee im totally confused

OpenStudy (anonymous):

I think you need to do a bit of studying. Try here http://www.purplemath.com/modules/lcm_gcf.htm Thing is, math builds up in layers, if you don't get a layer, you will just have trouble later on.

OpenStudy (anonymous):

true true true. thanks soo much for your help.:)

OpenStudy (anonymous):

You can try to keep posting your questions to see if someone will just give you the answer but most people will resist answering questions like the first three.

OpenStudy (anonymous):

Anyway, good luck.

OpenStudy (anonymous):

thanks.